OpenStudy (anonymous):

please help! will give medal When solving an equation in the form ax^2=c by taking square roots, how many solutions will there be?

OpenStudy (anonymous):

If we subtract c from both sides, we'd get: ax^2 - c =0. This has an x^2 term, which means this represents some sort of parabola. What do you know about parabolas? They have two roots, and cross the x-axis twice (at least the ones whose roots are not imaginary). So there will be two solutions to ax^2 = c.

OpenStudy (anonymous):

thank you! this makes a lot more sense now.

OpenStudy (anonymous):

That is a geometrical approach. We can instead divide 'a' on both sides and get: x^2 = c/a And then we just take the square root: x = sqrt[ c/a ]. However, where is the other answer? Well, the square root of a number is positive OR negative. For example, take 25. You either multiply (5)*(5) and get 25, or you can multiply (-5)*(-5) to get 25! So x = +sqrt[ c/a ] x = -sqrt[ c/a ]

OpenStudy (anonymous):

You're welcome!